40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 3 18:42:43 2014, Jenne, Update, LSC, Green PDH box boosts BoostCartoon.pdf
    Reply  Thu Sep 4 00:56:44 2014, Jenne, Update, LSC, Green PDH box boosts TFstage_newDesign_3Sept2014.pngBoostNoBoost_newDesign_3Sept2014.pngBoostOnly_newDesign_3Sept2014.pngD0901351-v1_3Sept2014.pdf
       Reply  Thu Sep 4 03:12:55 2014, ericq, Update, LSC, Green PDH box boosts boostedTF.pdfgain.jpg
       Reply  Thu Sep 4 16:45:10 2014, Jenne, Update, LSC, Green PDH box boosts IMG_1712.JPGIMG_1714.JPG
          Reply  Fri Sep 5 05:32:57 2014, Jenne, Update, LSC, Green PDH box boosts TRX_TRY_ALSvsIR_4Sept2014.pdf
             Reply  Fri Sep 5 21:13:57 2014, Jenne, Update, LSC, Green PDH out of loop ALS_XY_outOfLoop_5Sept2014.pdf
Message ID: 10462     Entry time: Fri Sep 5 21:13:57 2014     In reply to: 10458
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Green PDH out of loop 

I locked the arms with IR, and measured the beatnote spectra to get the out of loop noise for the PDH boxes. 

Unfortunately, we don't have a reference saved (that I can find), so we're going to have to compare to an elog of Koji's from a month ago.  I have created an out of loop ALS reference .xml file in the Templates/ALS folder.

ALS_XY_outOfLoop_5Sept2014.pdf

As we can see from Koji's elog 10302, the Xarm seems to have stayed the same, but the Yarm seems to have increased by about an order of magnitude below 100 Hz.  :(

ELOG V3.1.3-